EXAMPLE ONE -

Keep-or-Drop Decision

Petoskey Company produces three products: Alanson, Boyne, and Conway. A segmented income statement, with amounts given in thousands, follows:

Alanson Boyne Conway Total
Sales revenue $1,280 $185 $300 $1,765
Less: Variable expenses 1,115 45 210 1,370
Contribution margin $165 $140 $90 $395
Less direct fixed expenses:
Depreciation 50 15 10 75
Salaries 95 85 108 288
Segment margin $20 $40 $(28) $32

Direct fixed expenses consist of depreciation and plant supervisory salaries. All depreciation on the equipment is dedicated to the product lines. None of the equipment can be sold.

Assume that each of the three products has a different supervisor whose position would remain if the associated product were dropped.

Required:

CONCEPTUAL CONNECTION: Estimate the impact on profit that would result from dropping Conway. Enter amount in full, rather than in thousands. For example, "15000" rather than "15".

EXAMPLE 2-

Keep-or-Drop Decision

Petoskey Company produces three products: Alanson, Boyne, and Conway. A segmented income statement, with amounts given in thousands, follows:

Alanson Boyne Conway Total
Sales revenue $1,280 $185 $405 $1,870
Less: Variable expenses 1,115 45 304 1,464
Contribution margin $165 $140 $101 $406
Less direct fixed expenses:
Depreciation 50 15 10 75
Salaries 95 85 108 288
Segment margin $20 $40 $(17) $43

Direct fixed expenses consist of depreciation and plant supervisory salaries. All depreciation on the equipment is dedicated to the product lines. None of the equipment can be sold.

Assume that each of the three products has a different supervisor whose position would be eliminated if the associated product were dropped.

Required:

Conceptual Connection: Estimate the impact on profit that would result from dropping Conway. Enter amount in full, rather than in thousands. For example, "15000" rather than "15".

EXAMPLE 3 -

Keep-or-Drop Decision

Petoskey Company produces three products: Alanson, Boyne, and Conway. A segmented income statement, with amounts given in thousands, follows:

Alanson Boyne Conway Total
Sales revenue $1,280 $185 $300 $1,765
Less: Variable expenses 1,115 45 210 1,370
Contribution margin $165 $140 $90 $395
Less direct fixed expenses:
Depreciation 50 15 15 80
Salaries 95 85 116 296
Segment margin $20 $40 $(41) $19

Direct fixed expenses consist of depreciation and plant supervisory salaries. All depreciation on the equipment is dedicated to the product lines. None of the equipment can be sold.

Assume that, each of the three products has a different supervisor whose position would be eliminated if the associated product were dropped.

Assume that 20% of the Alanson customers choose to buy from Petoskey because it offers a full range of products, including Conway. If Conway were no longer available from Petoskey, these customers would go elsewhere to purchase Alanson.

Required:

Conceptual Connection: Estimate the impact on profit that would result from dropping Conway. Enter amount in full, rather than in thousands. For example, "15000" rather than "15".
Answer #1
1. Impact on profit that would result from dropping Conway (Supervisory Cost cannot be eliminated)
Particulars Alanson Boyne Total
Sales Revenue 1280000 185000 1465000
Less: Variable Expense 1115000 45000 1160000
Contribution Margin 165000 140000 305000
Less: Direct Fixed Expenses
           Depreciation 75000
           Salaries 288000
Net Profit -58000
Impact on Net Profit = -58000-32000 = -90000.00
It will result in Loss of $ 90,000.00
2. Impact on profit that would result from dropping Conway (Supervisory Cost can be eliminated)
Particulars Alanson Boyne Total
Sales Revenue 1280000 185000 1465000
Less: Variable Expense 1115000 45000 1160000
Contribution Margin 165000 140000 305000
Less: Direct Fixed Expenses
           Depreciation 75000
           Salaries 180000
Net Profit 50000
Impact on Net Profit = 50000-43000 = 7000.00
It will result in additional Profit of $ 7,000.00
2. Impact on profit that would result from dropping Conway (20% Alanson Also gone)
Particulars Alanson Boyne Total
Sales Revenue 1024000 185000 1209000
Less: Variable Expense 892000 45000 937000
Contribution Margin 132000 140000 272000
Less: Direct Fixed Expenses
           Depreciation 80000
           Salaries 180000
Net Profit 12000
Impact on Net Profit = 12000-19000 = -7000.00
It will result in Loss of $ 7,000.00